# Self-Questioning Language Models

### v3

@@ 1 Introduction
The post-training of large language models still relies heavily on hand-curated datasets (Yu et al., 2025a; Li et al., 2024), demanding substantial engineering effort and human supervision. In an attempt to alleviate this burden, researchers have developed unsupervised reward functions for reinforcement learning, which use proxies such as the model’s internal confidence (Zhao et al., 2025b; Prabhudesai et al., 2025) or the majority answer (Zuo et al., 2025) in the absence of ground-truth rewards or answers. However, these methods still presuppose the existence of well-formed input prompts or questions, and the bottleneck shifts from curating labeled answers to curating high-quality questions – a task that remains labor-intensive and not easily automated at scale. This highlights a critical gap in current methodologies: the lack of scalable, self-sustaining pipelines for generating meaningful questions and answers without human intervention.
@@ 2.2 Unsupervised Rewards for Reasoning
Ground-truth answers and external verifiers are often costly or impractical to obtain. This has motivated recent work exploring how far one can go without any form of external supervision. Notably, several studies have demonstrated that pre-trained language models are sufficiently well-calibrated to use their own confidence as a training signal. This has been operationalized through entropy minimization (i.e., reverse KL divergence to a uniform distribution) (Prabhudesai et al., 2025), and forward KL divergence (Zhao et al., 2025b). Another unsupervised reward strategy treats the majority prediction among sampled completions as the ”correct” answer (Zuo et al., 2025). These approaches rely solely on the model’s internal uncertainty, eliminating the need for labeled data or external evaluators. Building on this line of work, our approach goes a step further: not only does it not require ground-truth answers, it does not even require human-authored questions. Instead, the model itself is used to generate questions, making the entire pipeline fully self-supervised.
@@ 3.1 Reinforcement Learning for Language Models
Recently, several unsupervised reward functions $\mathcal{R}(x,y_{\text{pred}})$ have been proposed. These rewards do not rely on having access to the ground truth $y$, instead using majority voting (Zuo et al., 2025), entropy (reverse KL divergence) (Prabhudesai et al., 2025), forward KL divergence (Zhao et al., 2025b), or even random rewards (Shao et al., 2025).
@@ 4.2 Reward Functions
When the generator-verifier gap is small. In domains such as arithmetic, there is no easy way to verify solutions without performing reasoning that is similarly complex to what is required to generate the solution itself. For these cases, we do not ask the proposer to generate a solution, and instead use majority voting (Zuo et al., 2025) as a self-supervised reward for the solver. Specifically, for each problem, we sample $N$ generations from the model and use the majority answer as a proxy for the correct answer. All generations that match the majority answer are given a reward of 1 and all other generations are given a reward of 0.
@@ 5.1 Setups
Arithmetic. We prompt the proposer to generate a three-digit arithmetic problem and use the proposed problem as the input to the solver. The solver is trained via majority voting reward (Zuo et al., 2025) to reinforce the majority answer. The proposer reward is 1 if the problem is not too difficult (the majority answer only appears once, which means all the samples are different) and not too easy (the majority answer is equal to the number of samples $N$, which means all the samples were the same), and 0 otherwise. To evaluate the arithmetic ability of the model, we procedurally generate a test set of 4096 three-digit multiplication problems following the setup from TinyZero  [^3] https://github.com/Jiayi-Pan/TinyZero. We run our experiments with Qwen2.5-3B-Instruct.

### v4

@@ 1 Introduction
The post-training of large language models still relies heavily on hand-curated datasets (Yu et al., 2025a; Li et al., 2024), demanding substantial engineering effort and human supervision. In an attempt to alleviate this burden, researchers have developed unsupervised reward functions for reinforcement learning, which use proxies such as the model’s internal confidence (Zhao et al., 2025b; Prabhudesai et al., 2025) or the majority answer (Zuo et al., 2025; Shafayat et al., 2025) in the absence of ground-truth rewards or answers. However, these methods still presuppose the existence of well-formed input prompts or questions, and the bottleneck shifts from curating labeled answers to curating high-quality questions – a task that remains labor-intensive and not easily automated at scale. This highlights a critical gap in current methodologies: the lack of scalable, self-sustaining pipelines for generating meaningful questions and answers without human intervention.
@@ 2.2 Unsupervised Rewards for Reasoning
Ground-truth answers and external verifiers are often costly or impractical to obtain. This has motivated recent work exploring how far one can go without any form of external supervision. Notably, several studies have demonstrated that pre-trained language models are sufficiently well-calibrated to use their own confidence as a training signal. This has been operationalized through entropy minimization (i.e., reverse KL divergence to a uniform distribution) (Prabhudesai et al., 2025), and forward KL divergence (Zhao et al., 2025b). Another unsupervised reward strategy treats the majority prediction among sampled completions as the ”correct” answer (Zuo et al., 2025; Shafayat et al., 2025). These approaches rely solely on the model’s internal uncertainty, eliminating the need for labeled data or external evaluators. Building on this line of work, our approach goes a step further: not only does it not require ground-truth answers, it does not even require human-authored questions. Instead, the model itself is used to generate questions, making the pipeline fully self-supervised.
@@ 3.1 Reinforcement Learning for Language Models
Recently, several unsupervised reward functions $\mathcal{R}(x,y_{\text{pred}})$ have been proposed. These rewards do not rely on having access to the ground truth $y$, instead using majority voting (Zuo et al., 2025; Shafayat et al., 2025), entropy (reverse KL divergence) (Prabhudesai et al., 2025), forward KL divergence (Zhao et al., 2025b), or even random rewards (Shao et al., 2025).
@@ 4.2 Reward Functions
When the generator-verifier gap is small. In domains such as arithmetic, there is no easy way to verify solutions without performing reasoning that is similarly complex to what is required to generate the solution itself. For these cases, we do not ask the proposer to generate a solution, and instead use majority voting (Zuo et al., 2025; Shafayat et al., 2025) as a self-supervised reward for the solver. Specifically, for each problem, we sample $N$ generations from the model and use the majority answer as a proxy for the correct answer. All generations that match the majority answer are given a reward of 1 and all other generations are given a reward of 0.
@@ 5.1 Setups
Arithmetic. We prompt the proposer to generate a three-digit arithmetic problem and use the proposed problem as the input to the solver. The solver is trained via majority voting reward (Zuo et al., 2025; Shafayat et al., 2025) to reinforce the majority answer. The proposer reward is 1 if the problem is not too difficult (the majority answer only appears once, which means all the samples are different) and not too easy (the majority answer is equal to the number of samples $N$, which means all the samples were the same), and 0 otherwise. To evaluate the arithmetic ability of the model, we procedurally generate a test set of 4096 three-digit multiplication problems following the setup from TinyZero  [^3] https://github.com/Jiayi-Pan/TinyZero. We run our experiments with Qwen2.5-3B-Instruct.